I want to double check my work to see if I did it correctly. Please compare my answers with yours, thank you.

Based on the following information, calculate the sustainable growth rate (in percent, to two decimals places) for Aurora Co.: Profit margin-8.31 percent, capital intensity ratio = 0.59, debt-equity ratio = 0.51, net income $121944, dividends paid $52715. 0.101 043 0-1012213 10. Ian 2a 043224
Answer #1

Capital intensity ratio=Total assets/Sales

Hence Total assets=0.59Sales

Profit margin=Net income/Sales

Hence net income =0.0831Sales

Debt-equity ratio=Debt/equity

Hence debt=0.51equity

Total assets=debt+equity

=1.51equity

ROE=net income/equity

=(0.0831Sales)/equity

=(0.0831* Total assets/0.59)/(Total assets/1.51)

=0.212679661

Dividend payout ratio=Dividend/Net income

=(52715/121944)=0.432288591

Retention ratio=1-Dividend payout ratio

=1-0.432288591

=0.567711409

Sustainable growth rate=(ROE*Retention ratio)/[1-(ROE*Retention ratio)]

=(0.212679661*0.567711409)/[1-0.567711409]

=13.73%(Approx).